OpenStudy (jurynee):

Solve the equation for x. -2 - 3 4 x = 10 A) -16 B) 0 Eliminate C) 16 D) - 32 3

OpenStudy (jurynee):

D is a fraction

OpenStudy (baru):

is this the equation?\[-2 - \frac{3}{4}x=10\]

OpenStudy (jurynee):

the equation is supposed to be -2-3/4x=10

OpenStudy (baru):

can you do this step: add 2 to both sides of the equation

OpenStudy (jurynee):

-3/4x= 12

OpenStudy (jurynee):

?

OpenStudy (baru):

now multiply both sides by 4

OpenStudy (jurynee):

i do not know how to multiply a fraction

OpenStudy (baru):

to multiply fractions, just multiply numerator by numerator, and denominator by denominator ex:\[\frac{3}{7} \times \frac{11}{17} \\=\frac{3 \times 11}{7 \times 17}\\\]

OpenStudy (jurynee):

ok gtg bye

OpenStudy (baru):

bye :)

OpenStudy (qwertty123):

Nice job @baru

OpenStudy (jurynee):

it was -16 i got it right thank you so much @baru

OpenStudy (baru):

happy to help :)

OpenStudy (baru):

also, thank you guys for my 500th medal!! ^_^